OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (anonymous):

you really need to check all four and see which one works replace \(x\) by the first number, \(y\) by the second number and see if both \[y\leq x-5\] and \[y\leq -x-4\] are true btw there is no such thing as "less than or greater to" it is either "less than" \(<\) "less than or equal to" \(\leq\) "great than" \(>\) "greater than or equal to" \(\geq\) i guessed at the first. make sure not to say "less that or greater to" in class, or your teacher will think you are a fool

OpenStudy (anonymous):

we can check one if you like

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok but first which one is it? is it "less than or equal to"?

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

less than as in \(<\) or less than or equal to like this \(\leq \)

OpenStudy (anonymous):

the 1st 1

OpenStudy (anonymous):

ok lets try choice a)

OpenStudy (anonymous):

here \(x=1,y=10\) so the first inequality \(y<x-5\) would be come \(10<1-5\) which is certainly not true, so we don't even need to check the second one